### v3

# How to Correctly Report LLM-as-a-Judge Evaluations

Abstract
Large language models (LLMs) are widely used as scalable evaluators of model responses in lieu of human annotators. However, imperfect sensitivity and specificity of the LLM judges induce bias in naive evaluation scores. We propose a simple plug-in framework that corrects this bias and enables statistically principled uncertainty quantification. Our framework constructs confidence intervals that account for uncertainty from both the test dataset and a human-labeled calibration dataset. Additionally, it uses an adaptive strategy to allocate calibration samples for tighter intervals. Importantly, we characterize parameter regimes defined by the true evaluation score and the LLM judge’s sensitivity and specificity in which our LLM-based evaluation yields more reliable estimates than human-only evaluation. Moreover, we show that our framework remains unbiased under distribution shift between the test and calibration datasets, in contrast to existing approaches.
LLM, evaluation, LLM-as-a-judge, bias, correction, confidence interval, calibration, unbiased estimation, statistical inference
1 Introduction
The use of large language models (LLMs) as judges provides a cheap, scalable alternative to human evaluation for various tasks such as grading factual accuracy, assessing code quality or detecting harmful content (Zheng et al., 2023; Liu et al., 2023; Wang et al., 2023; Li et al., 2025; Gu et al., 2024). A common practice is to report the proportion of responses that an LLM judges as ‘correct’, denoted by $\hat{p}$. However, directly reporting raw LLM judgment scores is statistically problematic (Bross, 1954; Schwartz, 1985; Forman, 2005; Angelopoulos et al., 2023a; Boyeau et al., 2025; Fraser, 2024; Albinet, 2025). Because LLM judgments are inherently noisy, $\hat{p}$ generally deviates from the true accuracy and fails to provide a reliable estimate (Wang et al., 2024; Koo et al., 2024; Huang et al., 2025).
[…]
Figure 1: LLM judgments with error rates $1-q_{1}$ and $1-q_{0}$, where $q_{1}$ and $q_{0}$ are the LLM’s sensitivity and specificity, respectively.
To understand the source of the resulting bias in the judgment score $\hat{p}$, we first examine how an LLM judge makes errors when evaluating individual responses. As illustrated in Figure 1, an LLM may incorrectly label an ‘incorrect’ response as ‘correct’ or, conversely, mislabel a ‘correct’ response as ‘incorrect’. Let $q_{1}$ and $q_{0}$ denote the probabilities that the LLM correctly judges ‘correct’ and ‘incorrect’ responses, respectively. These quantities correspond to the sensitivity $q_{1}$ and specificity $q_{0}$ of the LLM judge.
(a) Raw LLM-based estimator $\hat{p}$
(b) Our bias-correction method
(c) Bias-corrected estimator $\hat{\theta}$
Figure 2: Bias and its adjustment in LLM-based judgment under imperfect LLM evaluators ($1-q_{0}=0.3$ and $1-q_{1}=0.1$). (a) When the true accuracy $\theta$ is low ($\theta<0.75$), the expected value of the raw LLM-based estimator $\mathbb{E}[\hat{p}$] overestimates $\theta$, whereas when $\theta$ is high ($\theta>0.75$), it underestimates $\theta$. (b) By accounting for the LLM judge’s sensitivity $q_{1}$ and specificity $q_{0}$, we obtain the bias-corrected estimator $\hat{\theta}$ and its confidence interval (CI). We estimate $q_{1}$ and $q_{0}$ using a calibration dataset with true labels from human evaluators paired with LLM judgments. Here, $n$, $m_{0}$, and $m_{1}$ denote the test-set size, the number of calibration samples whose true label is incorrect (e.g., $4+7=3$), and the number of calibration samples whose true label is correct (e.g., $4+7=11$), respectively. (c) The resulting estimator $\hat{\theta}$ is unbiased when the true values of $q_{0}$ and $q_{1}$ are known, or when a sufficiently large calibration dataset is available. A plug-in Python implementation of this procedure is provided at https://github.com/UW-Madison-Lee-Lab/LLM-judge-reporting.
As an illustration of how the raw LLM judgment scores can be biased, we consider the extreme scenario where $q_{1}=1$ and $q_{0}=0$. In this case, the LLM will judge every response as ‘correct’ and consequently, the average of the LLM’s judgments $\hat{p}$ will be equal to $1$ regardless of the true $\theta$.
In general, whenever the LLM is imperfect ($q_{0}+q_{1}<2$), the expected value of $\hat{p}$ deviates from the true accuracy $\theta$:
$\displaystyle\mathbb{E}[\hat{p}]=\theta+(2-q_{0}-q_{1})\Big(\frac{1-q_{0}}{2-q_{0}-q_{1}}-\theta\Big),$ (1)
implying positive bias at low $\theta$ and negative bias at high $\theta$ (see Section 5 for details). This is illustrated in Figure 2(a) for an LLM with $1-q_{0}=0.3$ and $1-q_{1}=0.1$. Specifically, $\mathbb{E}[\hat{p}$] overestimates $\theta$ when $\theta<0.75$ (blue line) and underestimates it when $\theta>0.75$ (red line), which arises from the two underlying judgment errors (green arrows). A high probability of incorrectly rejecting a ‘correct’ response (large $1-q_{1}$) induces negative bias at high accuracies, whereas a high probability of incorrectly accepting an ‘incorrect’ response (large $1-q_{0}$) induces positive bias at low accuracies.
This issue is not merely theoretical. With the increasing adoption of LLM-based evaluation, reported improvements may be driven by bias induced by judgment errors rather than true model gains. Since different evaluation procedures can induce biases of different magnitude and direction, some apparent advances in the literature may arise from evaluation artifacts rather than genuine improvements. This highlights the importance of careful comparisons and the use of calibrated judges when interpreting past findings, and motivates the need for a principled method for bias adjustment.
Fortunately, this bias can be corrected. When the sensitivity $q_{1}$ and specificity $q_{0}$ are known, a classical result (Rogan and Gladen, 1978) provides an exact adjustment. When they are unknown, they can be estimated from a calibration dataset with human-evaluated labels, and the resulting estimates $\hat{q}_{1}$ and $\hat{q}_{0}$ can be substituted into the correction formula. This leads to the bias-corrected estimator $\hat{\theta}$ in Figure 2(c).
More broadly, a variety of bias-correction methods have been studied for imperfect evaluators. In particular, prediction-powered inference (Angelopoulos et al., 2023a, b) provides a general framework that extends beyond categorical outcomes and can yield estimators of the true accuracy $\theta$ with lower variance; see Kloos et al. (2021); Chen et al. (2026). However, we show that existing methods can become biased under distribution shift between the test and calibration datasets, which can arise in LLM-as-a-judge settings (Jung et al., 2024). In contrast, the estimator $\hat{\theta}$ based on the method of Rogan and Gladen (1978) remains unbiased under such shifts. Accordingly, we adopt this estimator and focus our analysis on $\hat{\theta}$. Moreover, since the estimator $\hat{\theta}$ may exhibit higher variance than existing alternatives, we propose an adaptive strategy for allocating calibration samples that reduces its variance and, consequently, shortens the confidence interval for $\theta$ under our method.
In this paper, we provide a statistical framework for correctly reporting LLM-as-a-judge evaluations, as outlined in Figure 2(b). We summarize our key contributions below:
  • In Section 4, we introduce the bias-corrected estimator for the true accuracy $\theta$, derive the confidence interval, and further propose an adaptive strategy for allocating calibration samples that reduces its length. In Section 5, we establish theoretical guarantees for our method.
  • In Section 6, we characterize parameter regimes defined by $q_{1}$, $q_{0}$, and $\theta$ in which our estimator has smaller variance than that obtained from human-only evaluation.
  • In Section 7, we validate our method through extensive Monte Carlo simulations and real-world LLM-as-a-judge evaluations on the Chatbot Arena benchmark.
  • In Section 8, we show that estimators from existing bias-correction methods can still be biased under distribution shift between the test and calibration datasets, whereas our estimator remains unbiased under such shifts.
2 Related Work
Statistical Reliability in LLMs’ Evaluation.
Recent efforts have sought to formalize the statistical reliability of language model evaluations. While frameworks for calculating error bars and confidence intervals in benchmarks exist (Miller, 2024), they assume that LLM evaluators provide true labels. However, this assumption does not hold in the LLM-as-a-judge setting, where imperfect judges introduce bias into LLM-based scores. Recent works have examined these biases across various contexts, including natural language inference and test-time compute scaling (Godbole and Jia, 2025; Mukherjee et al., 2025; Feng et al., 2025). Aligned with this line of work, we propose a bias-corrected estimator together with statistically sound confidence intervals.
Estimation Methods to Mitigate Bias.
To address bias arising from imperfect evaluators (e.g., diagnostic or screening tests, or LLM-based judges), Rogan and Gladen (1978) proposed an adjustment that corrects prevalence estimates when sensitivity and specificity are known. Subsequent work has developed more general frameworks, most notably Prediction-Powered Inference (Angelopoulos et al., 2023a, b; Zrnic and Candès, 2024; Broska et al., 2025; Boyeau et al., 2025), which leverages a small set of true labels to improve estimation, as well as conditional calibration estimators (Buonaccorsi, 2010; Kloos et al., 2021; Meertens et al., 2022); see Section 8 for details. While these methods offer distinct advantages, such as variance reduction (Kloos et al., 2021; Chen et al., 2026), our approach builds on Rogan and Gladen (1978); Lang and Reiczigel (2014), adopting distributional assumptions that remain valid under shifts between the calibration and test datasets and are particularly well-suited to the LLM-as-a-judge setting.
3 Problem Setup: LLM-as-a-Judge
We consider the problem of evaluating responses based on human judgment. For example, each instance consists of a question and a corresponding response produced by a given model [^1] The model producing the response may be an LLM, but rule-based or statistical models are also possible.. We assume that humans can assess whether a response is ‘correct’ or ‘incorrect’. This assessment is formalized by a ground-truth labeling function $z:\mathcal{X}\to\{0,1\}$, where $z(x)=1$ indicates that humans judge the response in instance $x$ to be ‘correct’, and $z(x)=0$ otherwise. Applying the function to a random instance $X$ induces a binary random variable $Z=z(X)$.
Our goal is to estimate the true accuracy of the responses with respect to human judgment, defined as
$\theta:=\mathop{\rm Pr}\nolimits(Z=1)=\displaystyle\mathbb{E}[Z].$ (2)
Test Distribution and LLM-Based Judgment.
Let $\mathbb{P}$} denote the distribution over test instances to be evaluated. In practice, instead of relying on human annotators, an LLM is used as a surrogate judge. Let $\hat{Z}:=f_{\mathrm{LLM}}(X)\in\{0,1\}$ denote the LLM’s judgment, where $\hat{Z}=1$ indicates that the LLM marks the response as ‘correct’, and $\hat{Z}=0$ otherwise.
Let $n]:=\{1,\cdots,n\}$, where $n$ denotes the test-set size. Given a test dataset $\{x_{i}\}_{i\in[n]}$ sampled i.i.d. from $\mathbb{P}$}, the LLM produces predictions $\hat{z}_{i}:=f_{\mathrm{LLM}}(x_{i})$. The quantity reported in practice is the empirical fraction of instances judged as ‘correct’ by the LLM:
$\hat{p}:=\frac{1}{n}\sum_{i\in[n]}\hat{z}_{i}.$ (3)
This estimator targets the probability $p:=\mathop{\rm Pr}\nolimits(\hat{Z}=1)$ that the LLM judges a test instance as ‘correct’.
However, the LLM’s judgment $\hat{Z}$ does not necessarily coincide with the human-evaluated true label $Z$. We characterize the LLM judge using the following parameters:
$q_{1}:=\mathop{\rm Pr}\nolimits(\hat{Z}=1\!\mid\!Z=1),\;q_{0}:=\mathop{\rm Pr}\nolimits(\hat{Z}=0\!\mid\!Z=0).\;\;$ (4)
which correspond to the sensitivity (true positive rate) and specificity (true negative rate) of the LLM judge, respectively (Forman, 2008; Lang and Reiczigel, 2014). Consequently, the LLM may incorrectly reject responses that are truly ‘correct’ with probability $1-q_{1}$, or accept responses that are truly ‘incorrect’ with probability $1-q_{0}$. Unless the LLM is perfectly accurate (i.e., $q_{0}=q_{1}=1$), the naive estimator $\hat{p}$ in (3) is generally a biased estimator of $\theta$.
Calibration Distribution and Estimation of Sensitivity and Specificity.
Let $\mathbb{Q}$} denote the distribution over calibration instances. Unlike the test dataset, each calibration instance is evaluated by humans, so that both the true label $Z$ and the corresponding LLM judgment $\hat{Z}$ are observable.
Let $m$ denote the calibration-set size, and let $m_{1}$ and $m_{0}$ denote the numbers of calibration instances with $z_{j}=1$ and $z_{j}=0$, respectively, so that $m=m_{0}+m_{1}$. The index $j$ distinguishes calibration instances from test instances indexed by $i$. Using the calibration dataset, we estimate the sensitivity $q_{1}$ and specificity $q_{0}$ of the LLM judge in (4) as
$\hat{q}_{1}\!:=\!\!\sum_{j\in[m]}\!\frac{\mathbf{1}\{\hat{z}_{j}=1,z_{j}=1\}}{m_{1}},\;\hat{q}_{0}\!:=\!\!\sum_{j\in[m]}\!\frac{\mathbf{1}\{\hat{z}_{j}=0,z_{j}=0\}}{m_{0}},\!\!\!$
where $\mathbf{1}\{\cdot\}$ denotes the indicator function. In the main analysis, we assume $\mathbb{P}$$\mathbb{Q}$}. In Section 8, we relax this assumption and consider distribution shift, where $\mathbb{P}$$\neq{\mathbb{Q}}$.
Problem Statement.
Because the LLM judge is imperfect, the naive estimator $\hat{p}$ is generally biased for the true accuracy $\theta$ defined in (2), i.e., $\mathbb{E}[\hat{p}]\neq\theta$. Moreover, existing LLM-as-a-judge evaluations typically report this point estimate without quantifying uncertainty via confidence intervals.
Our objective is therefore twofold: (i) to construct a bias-corrected estimator of the true accuracy $\theta$, and (ii) to provide statistically sound confidence intervals that reflect uncertainty arising from both the test and calibration datasets.
4 Method to Correctly Report LLM-as-a-Judge Evaluations
In this section, we present a bias-corrected estimator $\hat{\theta}$ and a confidence interval for LLM-as-a-judge evaluations. We further propose an adaptive allocation strategy for constructing the calibration dataset that reduces the confidence interval length. Derivations and theoretical guarantees for the proposed method are provided in Section 5.
4.1 Mitigating Bias in Point Estimator
We begin with the setting in which the sensitivity $q_{1}$ and specificity $q_{0}$ in (4) are known. In this case, an unbiased estimator of the true accuracy $\theta$ in (2) is given by
$\hat{\theta}\,\big|\,q_{0},q_{1}=\frac{\hat{p}+q_{0}-1}{q_{0}+q_{1}-1}.$ (6)
In realistic settings, these parameters $q_{1}$ and $q_{0}$ are unknown and must be estimated from a calibration dataset. Substituting estimates $\hat{q}_{1}$ and $\hat{q}_{0}$ into (6) gives the bias-corrected estimator (Rogan and Gladen, 1978; Lang and Reiczigel, 2014):
$\boxed{\hat{\theta}=\frac{\hat{p}+\hat{q}_{0}-1}{\hat{q}_{0}+\hat{q}_{1}-1}}.$ (7)
4.2 Uncertainty Quantification via Confidence Interval
To quantify uncertainty in $\hat{\theta}$, we derive a confidence interval for $\theta$ that incorporates variance contributions from both the test and calibration dataset (Lang and Reiczigel, 2014):
$\boxed{\begin{aligned} &\tilde{\theta}+d\tilde{\theta}\\ &\pm z_{\alpha}\sqrt{\frac{\frac{\tilde{p}(1-\tilde{p})}{\tilde{n}}+\big(1-\tilde{\theta}\big)^{2}\!\cdot\!\frac{\tilde{q}_{0}(1-\tilde{q}_{0})}{\tilde{m}_{0}}+\tilde{\theta}^{2}\!\cdot\!\frac{\tilde{q}_{1}(1-\tilde{q}_{1})}{\tilde{m}_{1}}}{(\tilde{q}_{0}+\tilde{q}_{1}-1)^{2}}}\end{aligned}},\;\;$ (8)
where values outside the interval $[0,1]$ are truncated to 0 or $1$. Here, $z_{\alpha}$ denotes the $(1-\alpha/2)$ quantile of the standard normal distribution, e.g., $z_{0.05}=1.96$, and the adjusted quantities are defined as
$\tilde{n}=n+z_{\alpha}^{2},\quad\tilde{m}_{0}=m_{0}+2,\quad\tilde{m}_{1}=m_{1}+2,$ (9)
$\tilde{p}=\frac{n\cdot\hat{p}+z_{\alpha}^{2}/2}{n+z_{\alpha}^{2}},\quad\tilde{q}_{0}=\frac{m_{0}\cdot\hat{q}_{0}+1}{m_{0}+2},$ (10)
$\tilde{q}_{1}=\frac{m_{1}\cdot\hat{q}_{1}+1}{m_{1}+2},\quad\tilde{\theta}=\frac{\tilde{p}+\tilde{q}_{0}-1}{\tilde{q}_{0}+\tilde{q}_{1}-1},$ (11)
$d\tilde{\theta}=2z_{\alpha}^{2}\left(-(1-\tilde{\theta})\cdot\tfrac{\tilde{q}_{0}(1-\tilde{q}_{0})}{\tilde{m}_{0}}+\tilde{\theta}\cdot\tfrac{\tilde{q}_{1}(1-\tilde{q}_{1})}{\tilde{m}_{1}}\right).$ (12)
The confidence interval in (8) reflects uncertainty from both the test and calibration datasets through $\tilde{n}$, $\tilde{m}_{0}$, and $\tilde{m}_{1}$. As these sample sizes increase, the terms inside the square root decrease, resulting in a shorter confidence interval for $\theta$. Because LLM-as-a-judge evaluations can be run at scale with minimal cost, the test-set size $n$ can often be made extremely large. In the limit $n\to\infty$, test-set uncertainty vanishes, and the interval length is determined solely by the calibration sample sizes $m_{0}$ and $m_{1}$. This observation enables practitioners to target a desired interval length and determine the minimal calibration budget required to achieve it.
Figure 3 illustrates how the confidence-interval length decreases as the calibration dataset grows. The dashed curves correspond to calibration datasets with $m_{0}=m_{1}$. For example, when $\hat{p}=0.3$ is estimated from the test set and $\hat{q}_{0}=0.7$, $\hat{q}_{1}=0.9$ (red dashed curve), achieving an interval shorter than $0.1$ requires $m\approx 200$ calibration examples.
Figure 3: Confidence-interval length versus calibration-set size under $\hat{q}_{0}=0.7$, $\hat{q}_{1}=0.9$, and $n\to\infty$. We consider four test-set cases in which the naive estimator $\hat{p}$ takes one of the values in $\{0.3,0.5,0.7,0.9\}$. Dashed curves correspond to calibration sets with symmetric allocation across label types ($m_{0}=m_{1}$), while solid curves correspond to calibration sets using the adaptive allocation strategy in Algorithm 1, which results in shorter intervals.
Allocation Strategy to Reduce Confidence-Interval Length.
Furthermore, as the calibration dataset is collected independently, its label composition can sometimes be influenced through sampling strategies. As a result, asymmetric label sizes ($m_{0}\neq m_{1}$) are feasible. This flexibility matters because the two label types typically contribute asymmetrically to the overall uncertainty, so an imbalanced allocation can reduce the interval length.
Motivated by the benefits of asymmetric allocation, we introduce an adaptive allocation strategy in Algorithm 1. The algorithm first collects a small pilot calibration set (e.g., $m_{\mathrm{pilot}}=10$ per label type) to obtain preliminary estimates $\tilde{q}_{0}$ and $\tilde{q}_{1}$. It then computes the empirical error ratio $(1-\tilde{q}_{0})/(1-\tilde{q}_{1})$ and combines it with the test-set estimate $\hat{p}$ to determine an allocation $(m_{0},m_{1})$ that approximately minimizes the confidence-interval length in (8). As shown by the solid curves in Figure 3, this adaptive allocation gives shorter intervals under a fixed calibration budget than symmetric allocation. The optimality of this strategy in minimizing the confidence-interval length under a fixed calibration budget is established in the following section.
5 Theoretical Justification of Our Method
We provide theoretical justification for the method proposed in Section 4, including derivations, bias analysis, and optimality of the allocation strategy.
5.1 Mitigating Bias in Point Estimator
We compare the bias-corrected estimator $\hat{\theta}$ in (7) with the naive estimator $\hat{p}$ in (3). By the law of total probability,
$\displaystyle\mathbb{E}[\hat{p}]=p=(q_{0}+q_{1}-1)\cdot\theta+(1-q_{0}).$ (13)
Thus, $\mathbb{E}[\hat{p}]=\theta$ for all $\theta$ if and only if the LLM judge is perfect (i.e., $q_{0}=q_{1}=1$); otherwise, $\hat{p}$ is biased. When $q_{0}+q_{1}<2$, the expectation can be rewritten as
$\displaystyle\mathbb{E}[\hat{p}$] $=\theta+(2-q_{0}-q_{1})\Big(\frac{1-q_{0}}{2-q_{0}-q_{1}}-\theta\Big),$ (14)
which makes the direction of the bias explicit: when the true accuracy $\theta$ is smaller than the threshold $\tfrac{1-q_{0}}{2-q_{0}-q_{1}}$, the estimator $\hat{p}$ exhibits a positive bias, i.e., $\mathbb{E}[\hat{p}]>\theta$; conversely, when $\theta$ exceeds this threshold, the bias becomes negative.
Bias-Corrected Estimator.
Assuming $q_{0}+q_{1}>1$, inverting the above relation in (13) gives the estimator in (6). Replacing $q_{0}$ and $q_{1}$ with their empirical estimates $\hat{q}_{0}$ and $\hat{q}_{1}$ gives the bias-corrected estimator $\hat{\theta}$ in (7).
When $q_{0}$ and $q_{1}$ are known, $\hat{\theta}$ is unbiased. When they are estimated from a calibration dataset, $\hat{\theta}$ can exhibit bias. Nevertheless, the following result shows that $\hat{\theta}$ attains smaller bias than the naive estimator $\hat{p}$ in (3) as the calibration-set size grows. All proofs are provided in Appendix B.
Proposition 5.1.
Suppose that $m:=2m_{0}=2m_{1}$ and that $q:=q_{0}=q_{1}$ with $0.5<q\leq 1$. For sufficiently large $m\gtrsim 2q/(2q-1)^{2}$, the absolute bias of $\hat{\theta}$ in (2) is always smaller than that of $\hat{p}$ in (3) for all $\theta\in[0,1]$.
Even when $\hat{q}_{0}$ and $\hat{q}_{1}$ are estimated from data, $\hat{\theta}$ has smaller bias than $\hat{p}$ for sufficiently large calibration sets, with the required size depending on the LLM judge’s sensitivity and specificity: fewer samples suffice when the LLM has high sensitivity and specificity ($q\approx 1$), while substantially more are required as these approach chance level ($q\approx 0.5$).
5.2 Uncertainty Quantification via Confidence Interval
We quantify uncertainty in $\hat{\theta}$ arising from two sources: the test dataset used to estimate $p$ and the calibration dataset used to estimate $q_{0}$ and $q_{1}$. Applying the delta method and the binomial variance formulas for $\hat{p}$, $\hat{q}_{0}$, and $\hat{q}_{1}$, we obtain the asymptotic variance
$\displaystyle\mathrm{Var}\big(\hat{\theta}\big)\!=\!\frac{\frac{\hat{p}(1\!\!-\!\hat{p})}{n}+(1\!-\!\hat{\theta})^{2}\!\cdot\!\frac{\hat{q}_{0}(1-\hat{q}_{0})}{m_{0}}+\hat{\theta}^{2}\!\cdot\!\frac{\hat{q}_{1}(1-\hat{q}_{1})}{m_{1}}}{(\hat{q}_{0}+\hat{q}_{1}-1)^{2}}.\;\;\;\;\;$ (15)
A detailed derivation is provided in Appendix B.1.
Based on this variance, we construct a confidence interval for $\theta$ using the “add two successes and two failures” adjusted Wald approach (de Laplace, 1820; Agresti and Coull, 1998; Brown et al., 2001; Lang and Reiczigel, 2014). Specifically, we replace $\hat{p}$, $\hat{q}_{0}$, and $\hat{q}_{1}$ with their adjusted versions $\tilde{p}$, $\tilde{q}_{0}$, and $\tilde{q}_{1}$, as defined in (11). These adjustments can be interpreted as adding one (or $z_{\alpha}^{2}/2$) success and one (or $z_{\alpha}^{2}/2$) failure to each estimate, improving coverage accuracy for small sample sizes (Agresti and Caffo, 2000). Substituting the estimates gives the confidence interval in (8).
The adjustment also induces a small shift in the interval center (i.e., $d\tilde{\theta}$ in (12)) due to the dependence of $\hat{\theta}$ on $\hat{q}_{0}$ and $\hat{q}_{1}$. Its effect on interval length is negligible and therefore ignored, see Lang and Reiczigel (2014) for details.
Optimal Allocation for Minimizing Confidence-Interval Length.
We characterize the optimal allocation of calibration samples across label types under a fixed budget for minimizing the confidence-interval length. To this end, we define the error ratio $\kappa:=(1-\tilde{q}_{0})/(1-\tilde{q}_{1})$.
Proposition 5.2.
Suppose that $\tilde{q}_{0}$ and $\tilde{q}_{1}$ are close to $1$. Then the minimum length of the confidence interval defined in (8) is achieved when $\tilde{m}_{0}\approx(1/\tilde{p}-1)\sqrt{\kappa}\cdot\tilde{m}_{1}$.
This result provides guidance for allocating calibration samples between the two label types. In particular, when the LLM judge is less accurate at identifying ‘correct’ responses (i.e., when $\kappa$ is small) or when the naive estimator $\tilde{p}$, the proportion of test-set responses judged as ‘correct’ by the LLM, is large, more calibration samples should be allocated to estimating $\tilde{q}_{1}$ to minimize the interval length, resulting in an optimal allocation with $m_{0}<m_{1}$. In practice, we approximate the optimal allocation using Algorithm 1, which estimates the error ratio $\hat{\kappa}$ from a pilot calibration sample and allocates $m_{0}$ and $m_{1}$ according to Proposition 5.2.
6 When Are LLM-as-a-Judge Evaluations Preferable to Human-Only Evaluation?
In Section 5, we show that a human-labeled calibration dataset is essential for reliable LLM-as-a-judge evaluation because it enables bias correction. This raises a natural question: If human annotators are available, why not instead estimate the target accuracy $\theta$ directly from the human-only evaluation on the test set?
More concretely, suppose that a fixed evaluation budget of $m$ human annotations is available. One could either (i) use these $m$ annotations to construct a calibration dataset for estimating the LLM-judge’s sensitivity and specificity, and then apply bias correction to LLM-based evaluations, or (ii) apply them directly to $m$ test instances to estimate the true accuracy from human evaluation alone. We compare these two strategies and show that, for the same budget $m$, LLM-as-a-judge evaluation with our bias-correction method can be statistically more efficient than human-only evaluation in certain parameter regimes.
Parameter Regimes Where LLM-as-a-Judge Achieves Lower Variance.
Human-only evaluation relies on a limited annotation budget, which restricts how many test instances can be evaluated and therefore limits how much the variance of the estimator can be reduced. In contrast, LLM-as-a-judge evaluation can be applied to an arbitrarily large test set at negligible marginal cost, so that the dominant source of uncertainty stems from bias correction based on a finite calibration set. These differences reflect an inherent trade-off between a costly, near-perfect oracle that can be queried only on a small number of instances, versus an inexpensive but imperfect oracle that can be deployed at scale. Importantly, as the judge becomes more reliable (i.e., as $q_{0}$ and $q_{1}$ approach one), the calibration-induced uncertainty diminishes. Consequently, by aggregating large-scale LLM judgments with appropriate bias correction, it is possible to obtain an estimator whose variance is smaller than that of human-only evaluation.
It is worth clarifying why human-only evaluation can still exhibit variance even if the true label of every test instance is evaluated and therefore known. The quantity of interest is the true accuracy of the test distribution, i.e., the expected value of the true label indicating whether an instance is ‘correct’. In practice, we observe a random sample of $m$ instances from the test distribution, and different samples of size $m$ result in different empirical accuracies. Thus, the variance of the estimator arises from subsampling the test distribution, not from uncertainty in individual labels.
Figure 4: Comparison of variances between the LLM-as-a-judge estimator under our correction method, $\hat{\theta}$ in (11), and the human-only estimator, $\hat{\phi}$ in Proposition 6.1, when $m\to\infty$. Shaded regions indicate regimes where our LLM-as-a-judge evaluation is preferable to human-only evaluation in terms of estimator variance.
(a) Estimates
(b) Coverage probability
(c) Confidence interval length
Figure 5: Monte Carlo simulation for estimating $\theta$ under an imperfect LLM judge with $(q_{0},q_{1})=(0.7,0.9)$. We evaluate estimators across 21 values of $\theta\in[0,1]$, each visualized as a single point. Figure 5(a) reports the results from a single run, while Figure 5(b) and Figure 5(c) summarize averages computed over 10,000 replications. All experiments use a test dataset of size $n=1000$ and a calibration dataset of size $m=200$, and we use an equal allocation $m_{0}=m_{1}$ for Figure 5(a) and Figure 5(b). (a) The naive estimator $\hat{p}$ in (3) exhibits bias, while the bias-corrected estimator $\hat{\theta}$ in (7) closely recovers the true accuracy $\theta$ across all values. Shaded regions represent the 95% confidence intervals (CI). (b) Across all $\theta$, the coverage probability of the confidence interval remains consistently close to the nominal 95% level. (c) Given a fixed calibration budget of $m=200$, we compare two allocation strategies: an equal split ($m_{0}=m_{1}$) and the allocation proportional to $m_{0}\propto(1/\hat{p}-1)\sqrt{\kappa}\cdot m_{1}$ by using Algorithm 1. The proposed allocation gives shorter confidence intervals.
Figure 4 visualizes this comparison. The shaded regions indicate parameter regimes over $q_{0}$, $q_{1}$, and $\theta$ where the bias-corrected LLM-as-a-judge estimator has smaller variance than the human-only estimator. The variance advantage is most pronounced around $\theta=1/2$, where the intrinsic uncertainty of binary evaluation is largest, and the shaded region expands as the LLM judge becomes more accurate. The following proposition formalizes these parameter regimes.
Proposition 6.1.
[…]
$\!\!\!\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}\!\cdot\!\left(1+\frac{(1-\theta)\epsilon}{1-\theta-\epsilon}+\frac{\theta\epsilon}{\theta-\epsilon}\right)\!.\;$ (16)
Moreover, since $\epsilon=O(m^{-1/2})$, as $m\to\infty$ the sufficient condition in (16) reduces to the necessary and sufficient condition $\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}$, which is equivalent to
$\theta\in\left[\tfrac{1}{2}-\sqrt{\tfrac{1}{2}-\tfrac{1}{4(2q-1)^{2}}},\tfrac{1}{2}+\sqrt{\tfrac{1}{2}-\tfrac{1}{4(2q-1)^{2}}}\right].$ (17)
7 Empirical Validation of Our Method
We validate the theoretical results in Section 5 through Monte Carlo simulations and real-world benchmarks.
7.1 Monte Carlo Simulation
We evaluate the proposed method under the following parameter configuration. The LLM judge is characterized by $(q_{0},q_{1})=(0.7,0.9)$, and the true accuracy varies over $\theta\in\{0,0.05,0.10,\ldots,1\}$, yielding 21 settings. For each $(q_{0},q_{1},\theta)$, we generate a test dataset of size $n=1000$ and a calibration dataset of total size $m=200$, with equal allocation $m_{0}=m_{1}$ unless stated otherwise.
We report the naive estimator $\hat{p}$ in (3) and its confidence interval, as well as the bias-corrected estimator $\hat{\theta}$ in (7) with the confidence interval in (8). Each configuration is replicated 10,000 times to evaluate estimation accuracy, confidence-interval coverage, and interval length. Additional results are provided in Appendix D.
Results.
Figure 5(a) compares $\hat{p}$ and $\hat{\theta}$ in a single simulation run. The naive estimator $\hat{p}$ is biased, particularly overestimating the true accuracy $\theta$ when $\theta$ is small. In contrast, the bias-corrected estimator $\hat{\theta}$ closely matches the true accuracy across all values of $\theta$, consistent with Proposition 5.1.
Figure 5(b) reports the empirical coverage probability of the confidence interval based on $\hat{p}$ and the proposed confidence interval in (8). The interval based on $\hat{p}$ attains near-zero coverage except at a few values of $\theta$. In contrast, the proposed interval achieves coverage close to the nominal 95% level across all values of $\theta$. These results confirm that our confidence interval provides reliable uncertainty quantification.
To evaluate the benefit of optimal allocation, we compare symmetric allocation ($m_{0}=m_{1}=100$) with the allocation produced by Algorithm 1 under a fixed calibration budget of $m=200$. Algorithm 1 produces a calibration set that approximately follows the optimal allocation in Proposition 5.2. Figure 5(c) shows that Algorithm 1 yields shorter confidence intervals than symmetric allocation.
7.2 Chatbot Arena Benchmark
We evaluate our method on the Chatbot Arena benchmark (Chiang et al., 2024), where users vote on which of two anonymous models provides the better response to the same prompt. These pairwise comparisons are aggregated into each model’s win rate, defined as the probability that the target model’s response is preferred over its opponent’s.
We consider six target models studied in Zheng et al. (2023): Alpaca-13B (Taori et al., 2023), Claude-v1, FastChat-T5-3B, GPT-4, LLaMA-13B (Touvron et al., 2023), and Vicuna-13B (Chiang et al., 2023). For each target model, we construct response pairs consisting of one response generated by the target model and one response generated by a randomly selected opponent model. Chatbot Arena provides human preference labels for these pairs, which we treat as ground truth for computing the target model’s true win rate $\theta$. We then apply GPT-4.1-mini as an LLM judge to the same pairs to obtain LLM-judged preference labels, and compute the naive estimator $\hat{p}$ in (3) and our bias-corrected estimator $\hat{\theta}$ in (7).
For each trial, we randomly split the human-labeled pairs into a test set (90%) and a calibration set (10%). We estimate the judge’s sensitivity and specificity on the calibration set, and evaluate the bias of $\hat{p}$ and $\hat{\theta}$ on the held-out test set. We repeat this procedure 100 times with independent random splits and report results averaged across trials.
Figure 6: Average bias of winning rates for six models on Chatbot Arena, averaged over 100 random test (90%) / calibration (10%) splits. We compare the naive LLM-based estimator $\hat{p}$ in (3) with our bias-corrected estimator $\hat{\theta}$ in (7), using GPT-4.1-mini as the judge. The proposed method reduces bias for all models.
Table 1: Empirical coverage of the 95% confidence intervals in (8) and average interval lengths for Chatbot Arena win-rate estimates.
Model Alpaca Claude FastChat-T5 GPT-4 LLaMA Vicuna
Coverage 96% 95% 97% 97% 99% 99%
CI length 0.193 0.262 0.256 0.210 0.267 0.178
Table 2: Existing bias-correction estimators for $\mathbb{E}_{{\mathbb{P}}}[{\mathbf{Z}}]$. Here, $\mathbb{P}$} denotes the test distribution and $\mathbb{Q}$} the calibration distribution. Estimators (ii)–(iv) rely on the assumption that $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}({\mathbf{Z}})=\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}({\mathbf{Z}})$, whereas our estimator in (v) remains valid when $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}({\mathbf{Z}})\neq\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}({\mathbf{Z}})$.
Name Estimation formula Assumption
[…]
(ii) Calibration-only estimator $\mathbb{E}_{{\mathbb{Q}}}[{\mathbf{Z}}]$ $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}({\mathbf{Z}})=\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}({\mathbf{Z}})$
[…]
(v) Misclassification-adjusted estimator (Ours; $\hat{\theta}$ in (7)) $\big(\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})\big)^{\!-1}\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}$] $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})=\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})$
Results.
Figure 6 shows the average bias in win-rate estimation for each model on the Chatbot Arena benchmark. The naive estimator $\hat{p}$ exhibits non-negligible bias across all six models, whereas our estimator $\hat{\theta}$ reduces bias toward zero. Table 1 reports the coverage of the proposed 95% confidence intervals, which remains close to or slightly above the nominal level across models, indicating valid uncertainty quantification. Overall, our method provides a simple and practical bias-correction procedure for LLM-as-a-judge evaluation that requires only a small calibration dataset.
8 Robust Estimation under Distribution Shift
As mentioned in Section 1, a variety of bias-correction methods have been proposed for imperfect evaluators, typically under the assumption $\mathbb{P}$$\mathbb{Q}$} between the test distribution $\mathbb{P}$} and the calibration distribution $\mathbb{Q}$}. When this assumption is violated, bias correction can fail.
In practice, however, not every marginal or conditional probability under $\mathbb{P}$} need coincide with its counterpart under $\mathbb{Q}$}. Instead, only certain conditional distributions of the data-generating process may remain invariant across $\mathbb{P}$} and $\mathbb{Q}$}, while others may change. We therefore clarify which probabilistic assumptions are plausibly shared between $\mathbb{P}$} and $\mathbb{Q}$} in practice. Under minimal assumptions that are realistic in LLM-as-a-judge evaluation, we then characterize which methods remain unbiased under distribution shift.
Data-Generating Process for LLM-as-a-Judge Evaluation.
LLM-as-a-judge evaluation can be formalized through a data-generating process consisting of two probabilistic components: the marginal distribution of true labels $\mathop{\rm Pr}\nolimits(Z)$, which is determined by the dataset, and the conditional behavior of the LLM judge, $\mathop{\rm Pr}\nolimits(\hat{Z}\mid Z,\xi)$. Here, $\xi$ denotes auxiliary factors that may influence judgments, such as response length or stylistic attributes (Zhou et al., 2023; Dubois et al., 2024).
In this section, we only assume that the judge’s behavior depends on the true label and that it remains invariant across the test and calibration datasets:
$\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(\hat{Z}\mid Z)=\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{Z}\mid Z).$ (18)
This assumption is motivated by common practice in LLM-as-a-judge evaluation, where the same LLM judge and prompting strategy are used for both the test and calibration datasets. In Appendix 9, we discuss how this assumption can be relaxed to allow the judge behavior to depend on additional covariates $\xi$, namely $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(\hat{Z}\mid Z,\xi)=\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{Z}\mid Z,\xi)$.
In contrast, we consider the true label distributions to differ:
$\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z)\neq\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z).$ (19)
Such distribution shift can arise because datasets are collected in different ways: calibration datasets are constructed to analyze how the judge makes errors, whereas test datasets reflect realistic evaluation scenarios (Jung et al., 2024). Since calibration datasets require costly human annotation, they are collected in advance and remain fixed over time. By contrast, test data are often collected online or arrive in multiple batches, each potentially having a different underlying true accuracy. Consequently, the true label distribution of the test data can differ from that of the calibration data.
Under this setup, Bayes’ rule implies
$\mathop{\rm Pr}\nolimits(Z\mid\hat{Z})\propto\mathop{\rm Pr}\nolimits(\hat{Z}\mid Z)\mathop{\rm Pr}\nolimits(Z),$ (20)
[…]
Implicit Assumptions of Existing Estimators.
Following prior work (Kloos et al., 2021; Meertens et al., 2022), we compare five point estimators for estimating the true accuracy $\theta$ in (2) under the test distribution, that is, $\mathbb{E}_{{\mathbb{P}}}[Z]$. To facilitate comparison, we define
$𝐙^:=(Z^,1Z^),\displaystyle$$\mathbf$$^{\top},\qquad\hat{{\mathbf{Z}}}:=(\hat{Z},1-\hat{Z})^{\top}$, (21)
[…]
Effect of Distribution Shift on Estimator Bias.
Most existing estimators rely on invariance assumptions that fail under the distribution shift in true accuracy described in (19), even when the judge’s conditional behavior remains unchanged. In particular, the estimators in (ii)–(iv) require invariance of the marginal distribution $\mathop{\rm Pr}\nolimits(Z)$, which is violated when $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z)\neq\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z)$. In contrast, the misclassification-adjusted estimator in (v) depends only on the conditional distribution $\mathop{\rm Pr}\nolimits(\hat{Z}\mid Z)$ and does not rely on the marginal distribution $\mathop{\rm Pr}\nolimits(Z)$.
Following the Monte Carlo simulation in Section 7, we examine estimator behavior under distribution shift. We consider a setting where the test and calibration datasets share the same conditional judge behavior, while the calibration accuracy varies as $\theta_{{\mathbb{Q}}}\in[0.25,0.75]$, with the test accuracy held fixed at $\theta_{{\mathbb{P}}}=0.5$.
Figure 7 reports the resulting biases of each estimator. As established in previous sections, the naive estimator in (i) is biased in this setting. Estimators (ii)–(iv) also exhibit bias when $\theta_{{\mathbb{P}}}\neq\theta_{{\mathbb{Q}}}$, whereas the misclassification-adjusted estimator in (v), used in our method, remains unbiased. Additional simulation details are provided in Appendix E.
Figure 7: Effects of distribution shift $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z)\neq\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z)$ on estimator bias. We fix the true accuracy of the test distribution at $\theta_{{\mathbb{P}}}=0.5$ and vary the true accuracy of the calibration distribution as $\theta_{{\mathbb{Q}}}\in[0.25,0.75]$, corresponding to a distribution-shift ratio from $0.5$ to $1.5$. The misclassification-adjusted estimator in (v), used in our method, remains unbiased under this shift.
9 Conclusion
In LLM-as-a-judge evaluation, noisy judgments can induce bias in naive estimates. We thus introduce a bias-corrected estimator and provide statistically sound confidence intervals. To further reduce interval length, we show that calibration design using the proposed algorithm, which allocates samples across response types, can be an effective strategy.
Beyond bias correction, we show that when the LLM judge is sufficiently accurate, our framework can achieve lower variance than human-only evaluation. Moreover, under a mild and realistic assumption on the LLM judge, our estimator remains unbiased and robust to distribution shift between the test and calibration datasets. Several directions for future work are discussed in Appendix 9. We hope this work contributes to more reliable, interpretable, and transparent practices for LLM-based evaluation.
Future Work.
Several directions remain for future work. (1) Our method can be extended to account for auxiliary factors that influence LLM-as-a-judge evaluations beyond the true label $\mathbf{Z}$}. While our current analysis assumes that the LLM evaluation $\hat{{\mathbf{Z}}}$ depends only on the true label $\mathbf{Z}$}, in practice it may also be affected by additional nuisance factors $\xi$, such as response length or other stylistic attributes. This more realistic setting can be accommodated by allowing both the confusion matrix $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}},\xi)$ and the LLM evaluation $\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}\mid\xi]$ to depend on such factors. For example, one may estimate separate confusion matrices across different strata of $\xi$ (e.g., short vs. long responses) and perform bias adjustment within each group, followed by aggregation across strata. Formally, this corresponds to correcting $\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}\mid\xi]$ using $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}},\xi)$ at the stratum level and averaging over $\xi$. (2) The proposed method can be extended to a multinomial setting by generalizing $\big(\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})\big)^{-1}\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}$] in Section 8 to a multivariate formulation, where $\mathbf{Z}$} and $\hat{{\mathbf{Z}}}$ are represented as probability distributions over multiple response categories. Such an extension would increase the number of parameters, and additional structural assumptions, such as constraints on the confusion matrix $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})$, may be required to ensure stable estimation. (3) A conformal prediction framework could be incorporated to provide sample-specific uncertainty quantification by constructing statistically valid prediction regions around each point estimate.
Acknowledgements
This work was supported by the National Science Foundation (NSF) Award DMS-2023239, the NSF CAREER Award CCF-2339978, an Amazon Research Award, and a grant from FuriosaAI. In addition, it was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean Ministry of Science and ICT (MSIT) (RS-2024-00345351, RS-2024-00408003), by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by MSIT (RS-2023-00259934, RS-2025-02283048), and by the Brain Korea 21 program funded by the Korean Ministry of Education and the NRF (BK21 FOUR, No.5199990913980).
Impact Statement
This paper advances the field of machine learning by improving the statistical reliability of LLM-as-a-judge evaluations. By correcting bias from imperfect LLM evaluators and providing uncertainty quantification, our work helps reduce the risk of misleading performance claims and supports more rigorous evaluation practices. The contribution is methodological and does not introduce new models that affect users but instead strengthens existing evaluation pipelines. We view the broader societal impact as positive.
Appendix A Adaptive Strategy for Allocating Calibration Samples Across Labels
We propose an adaptive strategy for allocating calibration sample sizes across the two true-label types (‘correct’ and ‘incorrect’) using a small pilot calibration set. Leveraging pilot estimates of sensitivity $\tilde{q}_{1}$ and specificity $\tilde{q}_{0}$, Algorithm 1 attains an optimal allocation under a fixed total calibration budget, leading to shorter confidence intervals.
Algorithm 1 Adaptive strategy for allocating calibration sample sizes across true label types
Input: Total calibration budget $m$, pilot sample size $2m_{\mathrm{pilot}}$ with $2m_{\mathrm{pilot}}\leq m$, and the estimate $\hat{p}$ from the test dataset.
Output: Allocated calibration sample sizes $(m_{0},m_{1})$.
1:Pilot calibration.
2:   Collect $m_{\mathrm{pilot}}$ calibration examples with true label $z_{j}=0$, and $m_{\mathrm{pilot}}$ examples with $z_{j}=1$.
3:   Compute $\tilde{q}_{0}$ and $\tilde{q}_{1}$:
$\tilde{q}_{0}=\frac{\sum_{z_{j}=0}\mathbf{1}\{\hat{z}_{j}=0,z_{j}=0\}+1}{m_{\mathrm{pilot}}+2},\quad\tilde{q}_{1}=\frac{\sum_{z_{j}=1}\mathbf{1}\{\hat{z}_{j}=1,z_{j}=1\}+1}{m_{\mathrm{pilot}}+2}.$
4:   Compute the estimated error ratio:
$\hat{\kappa}=\frac{1-\tilde{q}_{0}}{1-\tilde{q}_{1}}.$
5:Compute adaptive allocation.
6:   Using the approximation in Proposition 5.2, compute the provisional allocation:
[…]
7:   Enforce pilot size:
$m_{1}=\min\big\{\max\{m_{1}^{\star},\,m_{\mathrm{pilot}}\},m-m_{\mathrm{pilot}}\big\},\quad m_{0}=m-m_{1}.$
Appendix B Proofs
B.1 Deriving the Variance of Estimators
Because $p$ follows a binomial distribution, the variance of $\hat{p}$ is
$\displaystyle\mathrm{Var}(\hat{p})=\hat{p}(1-\hat{p})/n.$ (22)
Similarly, we have $\mathrm{Var}(\hat{q}_{0})=\hat{q}_{0}(1-\hat{q}_{0})/m_{0}$ and $\mathrm{Var}(\hat{q}_{1})=\hat{q}_{1}(1-\hat{q}_{1})/m_{1}$.
We now derive the asymptotic variance of $\hat{\theta}$ using the delta method (Dorfman, 1938; Ver Hoef, 2012) for $\hat{\theta}=\frac{\hat{p}+\hat{q}_{0}-1}{\hat{q}_{0}+\hat{q}_{1}-1}$. The first order derivatives with respect to $\hat{p}$, $\hat{q}_{0}$, and $\hat{q}_{1}$ are
$\frac{\partial\hat{\theta}}{\partial\hat{p}}=\frac{1}{\hat{q}_{0}+\hat{q}_{1}-1},\qquad\frac{\partial\hat{\theta}}{\partial\hat{q}_{0}}=\frac{1-\hat{\theta}}{\hat{q}_{0}+\hat{q}_{1}-1},\qquad\frac{\partial\hat{\theta}}{\partial\hat{q}_{1}}=\frac{-\hat{\theta}}{\hat{q}_{0}+\hat{q}_{1}-1}.$ (23)
Assuming independence between the test dataset and the calibration dataset, the delta method gives
$\displaystyle\mathrm{Var}(\hat{\theta})=\frac{\hat{p}(1-\hat{p})/n+(1-\hat{\theta})^{2}\cdot\hat{q}_{0}(1-\hat{q}_{0})/m_{0}+\hat{\theta}^{2}\cdot\hat{q}_{1}(1-\hat{q}_{1})/m_{1}}{(\hat{q}_{0}+\hat{q}_{1}-1)^{2}}.$ (24)
B.2 Proofs of Propositions
Proposition B.1.
Suppose that $m:=2m_{0}=2m_{1}$ and that $q:=q_{0}=q_{1}$ with $0.5<q\leq 1$. For sufficiently large $m\gtrsim 2q/(2q-1)^{2}$, the absolute bias of $\hat{\theta}$ in (2) is always smaller than that of $\hat{p}$ in (3) for all $\theta\in[0,1]$.
Proof.
First, note that the bias of $\hat{p}$ in (3) is
$\displaystyle\mathbb{E}[\hat{p}]-\theta$ $=(q_{0}+q_{1}-1)\theta+(1-q_{0})-\theta=(2\theta-1)(1-q).$ (25)
Next, consider the bias of $\hat{\theta}$ in (2). By the second-order delta method, we have
$\displaystyle\mathbb{E}[\hat{\theta}$] $\approx\frac{p+q_{0}-1}{q_{0}+q_{1}-1}+\frac{1}{2}\left(-\frac{2(q_{1}-p)}{(q_{0}+q_{1}-1)^{3}}\cdot\frac{q_{0}(1-q_{0})}{m_{0}}+\frac{2(p+q_{0}-1)}{(q_{0}+q_{1}-1)^{3}}\cdot\frac{q_{1}(1-q_{1})}{m_{1}}\right)$ (26)
$=\theta-\frac{(q_{1}-p)}{(q_{0}+q_{1}-1)^{3}}\cdot\frac{q_{0}(1-q_{0})}{m_{0}}+\frac{(p+q_{0}-1)}{(q_{0}+q_{1}-1)^{3}}\cdot\frac{q_{1}(1-q_{1})}{m_{1}},$ (27)
which implies
$\displaystyle\mathbb{E}[\hat{\theta}]-\theta$ $\approx\frac{-(1-\theta)q_{0}(1-q_{0})/m_{0}+\theta q_{1}(1-q_{1})/m_{1}}{(q_{0}+q_{1}-1)^{2}}=\frac{1}{m}\cdot\frac{2q}{(2q-1)^{2}}\cdot(2\theta-1)(1-q).$ (28)
Hence, for sufficiently large $m$ satisfying $m\gtrsim 2q/(2q-1)^{2}$, we conclude the following for all $\theta\in[0,1]$:
$\big|\displaystyle\mathbb{E}[\hat{\theta}]-\theta\big|$ $\approx\left|\frac{1}{m}\cdot\frac{2q}{(2q-1)^{2}}\right|\cdot\left|(2\theta-1)(1-q)\right|<\left|(2\theta-1)(1-q)\right|=\big|\displaystyle\mathbb{E}[\hat{p}]-\theta\big|.$ (29)
Proposition B.2.
Suppose that $n\to\infty$ and $q:=q_{0}=q_{1}$ with $\frac{1}{2}+\frac{1}{2\sqrt{2}}<q\leq 1$. Let $M_{1}\sim\mathrm{Binomial}(m,\theta)$ denote the number of ‘correct’ responses by human evaluation, and define $\hat{\phi}:=M_{1}/m$. Let $\hat{\theta}$ be the bias-corrected estimator in (11). Fix $\delta\in(0,1)$ and define $\epsilon:=\sqrt{\tfrac{\log(2/\delta)}{2m}}$. If $\epsilon<\min\{\theta,1-\theta\}$, then with probability at least $1-\delta$,
$\displaystyle\mathrm{Var}(\hat{\theta})\leq\displaystyle\mathrm{Var}(\hat{\phi})\quad\text{whenever}\quad\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}\cdot\left(1+\frac{(1-\theta)\epsilon}{1-\theta-\epsilon}+\frac{\theta\epsilon}{\theta-\epsilon}\right).$ (30)
Moreover, since $\epsilon=O(m^{-1/2})$, as $m\to\infty$ the sufficient condition in (30) reduces to the necessary and sufficient condition $\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}$, which is equivalent to
$\theta\in\left[\frac{1}{2}-\sqrt{\frac{1}{2}-\frac{1}{4(2q-1)^{2}}},\frac{1}{2}+\sqrt{\frac{1}{2}-\frac{1}{4(2q-1)^{2}}}\right].$ (31)
Proof.
Since $M_{1}\sim\mathrm{Binomial}(m,\theta)$, we have
$\displaystyle\mathrm{Var}(\hat{\phi})=\frac{\theta(1-\theta)}{m}.$ (32)
From (15), as $n\to\infty$ the contribution of $\hat{p}$ vanishes, and hence
$\displaystyle\mathrm{Var}(\hat{\theta})=\frac{(1-\theta)^{2}\cdot\frac{q(1-q)}{m-M_{1}}+\theta^{2}\cdot\frac{q(1-q)}{M_{1}}}{(2q-1)^{2}}.$ (33)
Since $\hat{\phi}=M_{1}/m$, Hoeffding’s inequality implies that for any $\delta\in(0,1)$,
$\mathop{\rm Pr}\nolimits\Big(|\hat{\phi}-\theta|\geq\epsilon\Big)\leq\delta,\qquad\epsilon:=\sqrt{\frac{\log(2/\delta)}{2m}}.$ (34)
Hence, with probability at least $1-\delta$, we have $|\hat{\phi}-\theta|\leq\epsilon$. This implies $M_{1}\geq m(\theta-\epsilon)$ and $m-M_{1}\geq m(1-\theta-\epsilon)$, and therefore
$\frac{1}{M_{1}}\leq\frac{1}{m(\theta-\epsilon)},\qquad\frac{1}{m-M_{1}}\leq\frac{1}{m(1-\theta-\epsilon)}.$ (35)
Substituting these bounds into (33) gives
[…]
$=\frac{1}{m}\cdot\frac{q(1-q)}{(2q-1)^{2}}\left(\frac{(1-\theta)^{2}}{1-\theta-\epsilon}+\frac{\theta^{2}}{\theta-\epsilon}\right)$ (37)
$=\frac{1}{m}\cdot\frac{q(1-q)}{(2q-1)^{2}}\left(1+\frac{(1-\theta)\epsilon}{1-\theta-\epsilon}+\frac{\theta\epsilon}{\theta-\epsilon}\right).$ (38)
[…]
$\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}\cdot\left(1+\frac{(1-\theta)\epsilon}{1-\theta-\epsilon}+\frac{\theta\epsilon}{\theta-\epsilon}\right),$ (39)
which proves (16). Finally, as $m\to\infty$, we have $\epsilon=\sqrt{\log(2/\delta)/(2m)}\to 0$, and the above condition approaches $\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}$, which is then necessary and sufficient. Solving this quadratic inequality gives
$\theta\in\left[\frac{1}{2}-\sqrt{\frac{1}{2}-\frac{1}{4(2q-1)^{2}}},\frac{1}{2}+\sqrt{\frac{1}{2}-\frac{1}{4(2q-1)^{2}}}\right].$ (40)
The interval is real-valued if and only if $q\geq\frac{1}{2}+\frac{1}{2\sqrt{2}}$. ∎
B.3 Equivalence between the Misclassification-Adjusted Estimator and Our Estimator
We show that the misclassification-adjusted estimator $\left(\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})\right)^{-1}\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}$] in Section 8 is equivalent to our estimator $\hat{\theta}$ in (7). From the definition, we have
$\big(\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})\big)^{-1}\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}$] $=\begin{pmatrix}\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{Z}=1\mid Z=1)&\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{Z}=1\mid Z=0)\\ \mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{Z}=0\mid Z=1)&\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{Z}=0\mid Z=0)\end{pmatrix}^{-1}\begin{pmatrix}\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(\hat{Z}=1)\\ \mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(\hat{Z}=0)\end{pmatrix}$ (41)
$=\begin{pmatrix}\hat{q}_{1}&1-\hat{q}_{0}\\ 1-\hat{q}_{1}&\hat{q}_{0}\end{pmatrix}^{-1}\begin{pmatrix}\hat{p}\\ 1-\hat{p}\end{pmatrix}$ (42)
$=\frac{1}{\hat{q}_{0}+\hat{q}_{1}-1}\begin{pmatrix}\hat{q}_{0}&-(1-\hat{q}_{0})\\ -(1-\hat{q}_{1})&\hat{q}_{1}\end{pmatrix}\begin{pmatrix}\hat{p}\\ 1-\hat{p}\end{pmatrix}$, (43)
provided that $\hat{q}_{0}+\hat{q}_{1}\neq 1$. This gives
$\big(\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})\big)^{-1}\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}]=\frac{1}{\hat{q}_{0}+\hat{q}_{1}-1}\begin{pmatrix}\hat{p}+\hat{q}_{0}-1\\ \hat{q}_{1}-\hat{p}\end{pmatrix}$. (44)
In particular, the first component corresponds to the estimator of $\theta=\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z=1)$:
$\hat{\theta}=\frac{\hat{p}+\hat{q}_{0}-1}{\hat{q}_{0}+\hat{q}_{1}-1},$ (45)
which exactly matches $\hat{\theta}$ in (7).
Appendix C Code
All code used for this paper, including a plug-in Python implementation of the introduced method for LLM-as-a-judge evaluation, is available in https://github.com/UW-Madison-Lee-Lab/LLM-judge-reporting. To make this appendix self-contained, we provide below the key functions that compute the bias-corrected estimator and its confidence interval, corresponding to the method described in Section 4.
from math import sqrt
from scipy.stats import norm
def clip(x, low=0.0, high=1.0):
return max(low, min(high, x))
def point_estimator(p, q0, q1):
"""Compute the adjusted point estimate."""
th = (p+q0-1)/(q0+q1-1)
return clip(th)
def confidence_interval(p, q0, q1, n, m0, m1, alpha=0.05):
"""Compute the adjusted (1 - alpha) confidence interval."""
z = norm.ppf(1-alpha/2)
p, q0, q1 = (n*p+z**2/2)/(n+z**2), (m0*q0+1)/(m0+2), (m1*q1+1)/(m1+2)
n, m0, m1 = n+z**2, m0+2, m1+2
th = (p+q0-1)/(q0+q1-1)
dth = 2*z**2*(-(1-th)*q0*(1-q0)/m0+th*q1*(1-q1)/m1)
se = sqrt(p*(1-p)/n+(1-th)**2*q0*(1-q0)/m0+th**2*q1*(1-q1)/m1)/(q0+q1-1)
return clip(th+dth-z*se), clip(th+dth+z*se)
Figure 8: Python code implementation of the adjustment method described in Section 4 that computes the bias-corrected estimate and the $(1-\alpha)$ confidence interval for the true accuracy $\theta$. The inputs p, q0, and q1 are empirical estimates from the test and calibration datasets.
Appendix D Additional Results on Monte Carlo Simulation
To complement the main simulation results in Figure 5, we report additional Monte Carlo experiments across multiple configurations with test set size $n=1000$, calibration sizes $m\in\{200,500\}$, and $(q_{0},q_{1})\in\{(0.9,0.9),(0.7,0.7),(0.9,0.7),(0.7,0.9)\}$. All other aspects of the simulation design follow the main-text setup.
Across all $(n,m,q_{0},q_{1})$ configurations, the main simulation findings remain consistent: bias correction improves estimation accuracy, empirical coverage matches the nominal level, and optimized calibration allocation produces shorter confidence intervals. Each figure corresponds to a specific $(n,m,q_{0},q_{1})$ setting and contains three subplots.
D.1 Results for $n=1000$ and $m=200$
(a) Estimates
(b) Coverage probability
(c) Confidence interval length
Figure 9: Monte Carlo results for $(n,m,q_{0},q_{1})=(1000,200,0.7,0.7)$.
(a) Estimates
(b) Coverage probability
(c) Confidence interval length
Figure 10: Monte Carlo results for $(n,m,q_{0},q_{1})=(1000,200,0.7,0.9)$.
(a) Estimates
(b) Coverage probability
(c) Confidence interval length
Figure 11: Monte Carlo results for $(n,m,q_{0},q_{1})=(1000,200,0.9,0.7)$.
(a) Estimates
(b) Coverage probability
(c) Confidence interval length
Figure 12: Monte Carlo results for $(n,m,q_{0},q_{1})=(1000,200,0.9,0.9)$.
D.2 Results for $n=1000$ and $m=500$
(a) Estimates
(b) Coverage probability
(c) Confidence interval length
Figure 13: Monte Carlo results for $(n,m,q_{0},q_{1})=(1000,500,0.7,0.7)$.
(a) Estimates
(b) Coverage probability
(c) Confidence interval length
Figure 14: Monte Carlo results for $(n,m,q_{0},q_{1})=(1000,500,0.7,0.9)$.
(a) Estimates
(b) Coverage probability
(c) Confidence interval length
Figure 15: Monte Carlo results for $(n,m,q_{0},q_{1})=(1000,500,0.9,0.7)$.
(a) Estimates
(b) Coverage probability
(c) Confidence interval length
Figure 16: Monte Carlo results for $(n,m,q_{0},q_{1})=(1000,500,0.9,0.9)$.
Appendix E Experimental Setup for Distribution-Shift Analysis
We describe the experimental setup used to produce the distribution-shift results in Figure 7. The setup matches the Monte Carlo simulation in Section 7, except that we vary the true accuracy of the calibration dataset.
We fix the LLM judge behavior to $(q_{0},q_{1})=(0.7,0.9)$ and the test-set accuracy to $\theta_{{\mathbb{P}}}=0.5$, while varying the calibration-set accuracy over $\theta_{{\mathbb{Q}}}\in[0.25,0.75]$. For each setting $(\theta_{{\mathbb{P}}},\theta_{{\mathbb{Q}}})$, we generate a test dataset of size $1000$ with $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z=1)=\theta_{{\mathbb{P}}}$ and a calibration dataset of size $200$ with $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z=1)=\theta_{{\mathbb{Q}}}$. We average results over 10,000 Monte Carlo replications.